A pebble is hurdled upward from the top of a 1120-foot-tall building with an initial velocity of 144 feet per second. Neglecting air resistance, the height h(t) in feet of the pebble after t seconds is given by the polynomial function:
h(t)=-16t squared + 144t + 1120
When will the pebble hit the ground?

Answer by stanbon(75887)

A pebble is hurled upward from the top of a 1120-foot-tall building with an initial velocity of 144 feet per second. Neglecting air resistance, the height h(t) in feet of the pebble after t seconds is given by the polynomial function:
h(t)=-16t squared + 144t + 1120
When will the pebble hit the ground?
Ans: When h(t) = 0
Solve: -16t^2+144t+1120 = 0
Factor:
-16(t^2-9t-70) = 0
----
t^2-9t-70 = 0
Factor:
(t-14)(t+5) = 0
Ans: time = 14 seconds
===========================
Cheers,
Stan H.
============================